OpenStudy (anonymous):

What is the 8th term of the geometric sequence where a1 = 243 and a4 = 9?

OpenStudy (anonymous):

Find the common ratio r first using the explicit def of geo seq\[a(n)=a _{1}r^{n-1}\]\[9=243r^{4-1}\]\[r^3=27\]\[r=3\]

OpenStudy (anonymous):

r should equal 1/3 tho...

OpenStudy (anonymous):

as 243*3^2=6561...

OpenStudy (anonymous):

or your initial equation should have been a1/r^(n-1)

OpenStudy (anonymous):

other than that, ya

OpenStudy (anonymous):

correct r=1/3 \[a(8)=243\left( \frac{1}{3} \right)^{8-1}=\frac{1}{9}\]

OpenStudy (anonymous):

these are the answers given to choose from 0.0123456 0.037037 0.11111 0.3333

OpenStudy (anonymous):

1/9=0.1111111

OpenStudy (anonymous):

thanx champ

OpenStudy (anonymous):

What is the simplified form of